OpenStudy (aravindg):

let f(x)=e^x and g(x)=log x find a (f+g)(1) b (fg)(1) c (fog)(1) d (2f)(1)

OpenStudy (lalaly):

-.-

OpenStudy (aravindg):

??

OpenStudy (aravindg):

answer plzz

OpenStudy (anonymous):

answerrrr plzzzzzzzz you can't even type please

OpenStudy (aravindg):

please

OpenStudy (anonymous):

\[(F + G)(1) = e + 0 = e\] \[(FG)(1) = e*0 = 0\] \[(F\ o\ G)(1) = e^0 = 1\] \[2F(1) = 2e\]